OpenStudy (anonymous):

Multiplying a number by which of the following would result in a smallest number? 10^-1 10^0 10^1 10^2

OpenStudy (anonymous):

its confusing because its a trick question and i dont know which it would be

OpenStudy (anonymous):

10^0 is 1

OpenStudy (anonymous):

10 ^1 is 10

OpenStudy (amilapsn):

10^-1=0.1 10^0=1 10^1=10 10^2=100

OpenStudy (anonymous):

10^ 2 is 100

OpenStudy (anonymous):

so its a =)

OpenStudy (anonymous):

think of a dialation on a graph =)

OpenStudy (anonymous):

how is 10^0=1 ?

OpenStudy (anonymous):

wouldn't it be just 10?

OpenStudy (amilapsn):

10^0=10^(1-1)=10^1/10^1=1

OpenStudy (anonymous):

Anything to the zero power is 1

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

sooooooooooo

OpenStudy (anonymous):

do you get it or do you not and thats why the questions still open ?